An accounting intern for a local CPA firm was reviewing the financial statements of a client in the electronics industry. The intern noticed that the client used the FIFO method of determining ending inventory and cost of goods sold. When she asked a colleague why the firm used FIFO instead of LIFO, she was told that the client used FIFO to minimize its income tax liability. This response puzzled the intern because she thought that LIFO would minimize income tax liability. Required: What would you tell the intern to resolve the confusion?
Answer #1

Let us understand this with the following example to clear all the confusion

Case 1

For the month of March

Type Date Qty Rate Value
Opening 1/3/XX 100 5 500
Purchase 4/3/XX 40 4 160
Purchase 11/3/XX 60 3.5 210
Sale 12/3/XX 150 7 1050
Purchase 19/3/XX 120 3 360
Sale 27/3/XX 100 7 700
Ending 31/3/XX 70

Calculate ending Inventory and COGS as per both LIFO and FIFO

1. LIFO

Ending Inventory = 20 X 3 + 50 X 5 = 310
COGS = Op + Purchases - Cl = 500 + 730 - 310 = 920

2. FIFO

Ending Inventory = 70 X 3 = 210
COGS = 500 + 730 - 210 = 1020

As can be seen clearly, COGS is higher under FIFO. Higher COGS mean LOWER profits.

So, the intern's confusion stems from only one situation which will be the reverse of the above situation. As can be seen in the above table that PURCHASE PRICES are falling.

If the Purchase prices were RISING then LIFO would have given LOWER PROFITS to reduce the tax liability. But if it is other way then FIFO can give lower profits.